# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v3

@@ II.2.1 Mixing of Electron Densities
Besides the linear mixing method, ABACUS implements Pulay’s direct inversion in iterative subspace (DIIS) method Pulay (1980, 1982) and Broyden’s quasi-Newton approach Johnson (1988). Fig. 3 illustrates the electron density mixing strategies implemented in ABACUS to guarantee stable SCF convergence.
Magnetic calculations usually exhibit significantly poorer convergence behavior compared to their non-magnetic counterparts, frequently encountering two distinct failure cases, i.e., divergence of SCF or convergence to excited magnetic configurations. These issues are due to the complex coupling between charge and magnetization degrees of freedom when solving the Kohn-Sham equations. To address these challenges, ABACUS implements a mixing scheme that explicitly accounts for spin polarization. The residual vector $R$ is reformulated to simultaneously track both charge and magnetization density variations
@@
where $n$ determines the history length, $p_{j}$ are Pulay/Broyden coefficients, and the $q$-dependent terms represents a modified Kerker preconditioner Zhou et al. (2018). The damping parameters $q_{\rho}$ and $q_{m}$ control the suppression of long-wavelength oscillations in charge and magnetization densities, respectively.
ABACUS provides several choices to optimize magnetic calculations, beginning with the implementation of separate preconditioning parameters for charge $q_{\rho}$ and magnetization $q_{m}$ densities to address their distinct physical characteristics and convergence behaviors. By default, the code deactivates magnetic density preconditioning (setting $q_{m}$ = 0) to prioritize numerical stability in most cases, while still providing users the flexibility to enable magnetic preconditioning when enhanced convergence for challenging magnetic systems is required. This approach combines robust default settings with customizable options, allowing the code to automatically handle routine cases while permitting users to fine-tune the mixing parameters for complex magnetic configurations. We compared the magnetic convergence for several magnetic materials using two versions of the ABACUS (v3.4 and v3.5). All computational details are available online mag (2024). As illustrated in Fig. 4(a), a marked difference in SCF convergence behavior is observed between the two versions. Our findings demonstrate that ABACUS v3.5 achieves significantly enhanced convergence efficiency when compared to the version 3.4 of ABACUS.
@@
Figure 4: (a) Differences in SCF performance score (vide infra) between v3.4 and v3.5 on the selection set of 21 examples. Prior to v3.4, the residual was defined by $\rho_{\uparrow}$ and $\rho_{\downarrow}$, whereas starting from v3.5, the residual is defined by the charge density and the magnetic density, and they are mixed separately. Since some test cases failed to converge in v3.4, we present results with convergence scores rather than directly using convergence steps. The max number of iterations is set to 100 in all calculations, and the SCF convergence threshold scf_thr=1e-6. The SCF score is defined by $|\rm log_{10}(\delta\rho_{\rm last})|\times 10\times\frac{100}{\rm N_{step}}$, where $\delta\rho_{\rm last}$ is the density difference $R(\rho_{in})$ of the last iteration and $\rm N_{step}$ is the number of convergence iterations. (b) Difference in SCF convergence step ($\rm N_{step}$) for the DFT+U calculation with only mixing charge density and both mixing charge density and density matrix. Here, “JT” represents a structure under Jahn–Teller distortion Halcrow (2013). For the sake of reproduction, we have made public all the details of the calculations in the link mag (2024), where one can find the complete report and download all input files.
@@
The charge mixing scheme in DFT+U calculations can be further enhanced by an automated U-Ramping procedure that gradually increases the U parameter during SCF iterations Meredig et al. (2010); Dorado et al. (2013). We performed DFT+U calculations on a series of selected magnetic materials (source files available online mag (2024)) using ABACUS v3.6, with particular focus on comparing the SCF convergence behavior between calculations with and without density matrix mixing. Fig. 4(b) demonstrates the reduction of SCF convergence steps after the density matrix mixing method is adopted.
@@ II.2.2 Smearing Methods
In metallic systems, SCF iterations frequently encounter convergence difficulties stemming from charge density oscillations. These oscillations arise because the electron occupation number drops abruptly to zero at the Fermi energy, with electronic states near the Fermi surface frequently switch between filled and empty electronic states. Furthermore, there exists no variational principle that regulates the convergence of the system’s total energy with resepct to $k$-point sampling. Smearing methods mitigate this issue by introducing a controlled degree of partial occupation for Kohn-Sham orbitals in the vicinity of the Fermi energy, effectively damping numerical instabilities and promoting convergence. A few smearing methods that have been implemented in ABACUS are listed below.
@@
Second, the Fermi-Dirac smearing scheme is intrinsically aligned with finite-temperature density functional theory Mermin (1965), where the electronic temperature $T$ governs the statistical electron distribution. Crucially, within this framework the smearing width ($k_{B}T$) arises as a natural consequence of thermal excitations in the electronic system, rather than serving as an empirically tuned broadening parameter. The formula is
@@
where $k_{B}$ denotes Boltzmann’s constant. However, in comparison to Gaussian smearing, the drawback of the Fermi-Dirac smearing method lies in its long tail, which necessitates the calculation and storage of a large number of unoccupied states to obtain the charge density.
@@
Fourth, to address possible negative occupancies that may lead to negative total electron density during SCF and numerical instability in the MP method, Marzari et al. introduced the cold-smearing method Marzari et al. (1999), where the broadening function takes the form of
@@ II.3 Geometry Relaxation
This minimization problem can be solved using various optimization algorithms, including the steepest descent (SD), the conjugate gradient (CG), the Broyden-Fletcher-Goldfarb-Shanno (BFGS) Head and Zerner (1985), and the Fast Inertial Relaxation Engine (FIRE) Bitzek et al. (2006) methods. At each iteration step, the atomic positions are updated according to
@@
where $\Omega$ denotes the cell volume and $\epsilon_{\alpha\beta}$ is the strain tensor components. Next, different geometry relaxation strategies can be employed. For example, one approach to perform cell relaxation consists of two sequential optimization stages. First, optimizing atomic positions with lattice vectors held fixed, followed by adjusting lattice vectors while keeping ionic positions constant. An alternative strategy involves optimization of both atomic positions and lattice vectors at the same time. This simultaneous refinement allows for dynamic coupling between atomic rearrangements and lattice deformation, potentially accelerating convergence to the equilibrium structure. The geometry optimization methods implemented in ABACUS have been applied to study alloys Liu et al. (2024a), interfaces Liu, Liu, and Chen (2021), slabs Liu et al. (2022), and low-dimensional materials Chen et al. (2022b); Liu and Chen (2025), etc.
@@ II.4 Molecular Dynamics
Central to MD simulations is the potential energy surface (PES), which describes atomic interactions and can be derived through various methods. Quantum mechanical approaches like density functional theory (DFT) can provide high-accuracy PES calculations but are computationally intensive. Alternatively, machine learning potentials Wang et al. (2018); Zeng et al. (2023) offer a balance between accuracy and efficiency by learning the PES from DFT training data, enabling simulations of larger systems or longer timescales. Although classical MD methods enable rapid PES evaluations for systems, their reliance on empirical force fields (such as Lennard-Jones potential Jones and Chapman (1924a, b) results in lower accuracy compared to the other two approaches.
Once the PES is established, atomic forces are computed via gradients of the total energy with respect to atomic positions, which are then integrated over time to predict atomic trajectories. The velocity Verlet algorithm Swope et al. (1982), a widely used integration scheme that ensures long-term stability, updates atomic positions $\mathbf{r}(t+\Delta t)$ and velocities $\mathbf{v}(t+\Delta t)$ in discrete time steps $\Delta t$.
MD simulations can be performed under various statistical ensembles, each imposing different constraints to model specific experimental conditions. ABACUS supports the following ensembles. First, the NVE (microcanonical) ensemble conserves particle number $N$, volume $V$, and total energy $E$, simulating isolated systems where energy exchange with the environment is negligible. Second, for simulations requiring constant temperature $T$, the NVT (canonical) ensemble Glenn J. Martyna and Klein (1996) employs thermostats such as velocity rescaling, a straightforward but non-equilibrium approach, or the Nosé-Hoover chain method Nosé (1984); Hoover (1985); Martyna, Klein, and Tuckerman (1992), which rigorously incorporates auxiliary degrees of freedom to maintain temperature via Hamiltonian dynamics. In addition, the Langevin dynamics Schneider and Stoll (1978) couples the system to a heat bath through stochastic and frictional forces, ensuring that the system reaches and maintains the desired temperature over time, mimicking conditions where the system exchanges energy with its environment. Third, the NPT (isothermal-isobaric) ensemble maintains constant pressure $P$ and temperature $T$ using barostats like Parrinello-Rahman.
Besides the above methods, ABACUS also supports the multi-scale shock technique (MSST) Reed, Fried, and Joannopoulos (2003), enabling the simulation of shock wave propagation in materials, thereby extending the range of accessible timescales for investigating shock phenomena within such non-equilibrium frameworks.
@@
Figure 6: Mean square displacement (MSD) as a function of molecular dynamics simulation time (in fs) for the three elements in the stage-4 AlCl4 graphite intercalation compound (GIC) using ABACUS and FHI-AIMS packages. (Adapted with permission from Phys. Rev. Applied 12, 044060 (2019). Copyright 2019 American Physical Society.)
Wang et al. Wang et al. (2019b) studied the intercalation behavior of AlCl4 molecule in graphite by employing BOMD simulations. Fig. 6 illustrates the mean-square displacements (MSD) as simulated using FHI-AIMS and ABACUS packages. The results show rapid diffusion within the graphite layers, evident from the steep rises in their MSD. Additionally, the MSD trajectories for the Cl atoms feature periodic oscillations that are ascribed to the rotational motion of these atoms around an axis orthogonal to the graphene layers during intercalation. This study revealed that AlCl4 molecule retains its tetrahedral structure post-intercalation into graphite. Based on diffusion behavior, they suggest that an aluminum chloride graphene battery with AlCl4 could deliver a higher cathode specific capacity, elevated voltage, and improved rate capability.
Figure 7: Radial distribution functions (RDFs) $g(r)$ for B systems at a density of 2.46 $\mathrm{g/cm^{3}}$ and the temperatures are (a) 86 eV and (b) 350 eV. Some of the $g(r)$ are derived from Ext-FPMD Blanchet et al. (2022a), as well as the sDFT calculations employed the PBE Perdew, Burke, and Ernzerhof (1996) and corrKSDT Karasiev, Dufty, and Trickey (2018) XC functionals. The number of B atoms is set to 32 in AIMD simulations. In addition, DPMD denotes the model trained by the traditional DP method Wang et al. (2018), whereas DPMD-T indicates the TDDP method applied model training as reported by Zhang et al. Zhang et al. (2020b). NN is the number of B atoms in a cell. (Adapted with permission from Matter. Radiat. Extremes 9, 015604 (2024). Copyright 2024 AIP Publishing.)
Chen et al. Chen et al. (2024a) combined stochastic density functional theory (sDFT) with BOMD to explore warm dense matter systems at temperatures ranging from several tens of eV to 1000 eV. They also trained machine-learning-based interatomic models using the first-principles data and employed these models to examine large systems via BOMD simulations. Furthermore, they evaluated the structural and dynamic characteristics, as well as the transport coefficients, of warm dense matter. Fig. 7 shows the radial distribution functions $g(r)$ of warm dense B with a density of 2.46 $\mathrm{g/cm^{3}}$ at 86 and 350 eV, and the sDFT results are in excellent agreement with those obtained from extended first-principles molecular dynamics (Ext-FPMD) Zhang et al. (2016); Blanchet et al. (2022b).
Ma et al. Ma et al. (2024) combine AIMD and finite-temperature orbital-free DFT (FT-OFDFT) with a nonlocal free energy functional XWM to study a variety of warm dense matter systems such as the Si, Al, H, He, and H-He mixtures. The KS-DFT calculations were performed using ABACUS, while the OF-DFT calculations were carried out with ATLAS Mi et al. (2016). The XWMF functional is expected to be a good choice for the realistic simulations of warm dense matter systems covering a broad range of temperatures and pressures.
@@ II.5 Implicit Solvation Model
We place the solute in a cavity surrounded by a continuum dielectric medium characterized by the relative permittivity of the solvent, as employed in a previous work. Mathew et al. (2014). We describe the dielectric response in terms of the solute’s electron density, considering the polarization of solvents in response to the electronic structures of solute, the effects of cavitation and dispersion, and the reaction of solute system to the presence of the solvent.
@@
We benchmark the accuracy of the implicit solvation implementation by calculating molecular solvation energies $E_{\mathrm{sol}}$, which is defined as the total energy difference between a solvated condition and a vacuum condition in ABACUS and comparing them against VASPsol-calculated values Mathew et al. (2019). We can see from Fig. 9 that the solvation energies from ABACUS agree well with those produced by the VASPsol package. The minor discrepancies between the solvation energies computed by the two methods may be attributed to the pseudopotential difference. Note that we use projector-augmented-wave (PAW) potentials in VASPsol and the norm-conserving pseudopotentials in ABACUS.
@@
ABACUS also supports adding dipole correction and compensating charge when modeling surfaces, which were utilized in a recent work Sun et al. (2025). The periodic boundary conditions imposed on the electrostatic potential create an artificial electric field across a slab. By introducing an isolated slab-shaped density distribution $\rho(\mathbf{r})$ that is normal to the $z$-axis, a dipole correction Bengtsson (1999) is added to the bare ionic potential to compensate for the artificial dipole field within the context of periodic supercell calculations.
@@
Modeling a constant-potential electrochemical surface reaction requires the adjustment of electron numbers in a simulation cell. Simultaneously, we must preserve the supercell’s neutrality under the periodic boundary conditions. Thus, a distribution of compensating charge needs to be implemented in the vacuum region of surface models when extra electrons are added to or extracted from the system. The compensating charge implemented in ABACUS follows the methodology developed by Brumme et al. Brumme, Calandra, and Mauri (2014). We assume that the monopole with a total charge of $n_{\mathrm{dop}}$ per unit cell is located at $z_{\mathrm{mono}}$ along the $z$ axis. In this case, the effective potential $V_{\mathrm{mono}}(\mathbf{r})$ is added to the Hamiltonian, and the additional term $E_{\mathrm{mono}}$ is included in the total energy. Since $V_{\mathrm{mono}}(\mathbf{r})$ is independent of the electron density, it is unnecessary to update in a self-consistent manner. We also implemented the correction on the ionic forces induced by the presence of the monopole, allowing for calculating the electronic structure and complete structural relaxation in the field-effect configuration.
@@
Figure 10: Electrostatic energy of an electron along the $z$ direction perpendicular to the Pt surface calculated by Quantum ESPRESSO and ABACUS Sun et al. (2025). The gray vertical dashed line represents the position of the compensating charge plate, while the blue dashed line corresponds to the dipole correction. A schematic structure plot of the testing interface model is aligned below the average electrostatic energy curves. Gray, red, and white spheres correspond to platinum, oxygen, and hydrogen atoms, respectively.
@@
We output the electrostatic energy of an electron along the $z$ direction perpendicular to the Pt surface for a testing system with an added electron number $N_{e}^{\mathrm{extra}}=0.2$ in the simulation cell. The testing system is a (3×3) Pt (111) surface slab composed of three atomic layers, as illustrated in Fig. 10. The modeled electrode surface contains 27 Pt atoms, with a monolayer of hydrogen coverage and six explicit water molecules. We place a compensating charge plate in the vacuum region above the water layer, and a dipole correction is also included in DFT calculations. Fig. 10 shows that the electrostatic energy results from ABACUS are in perfect agreement with those computed from Quantum ESPRESSO Giannozzi et al. (2020).

### v4

@@ II.2.1 Mixing of Electron Densities
Besides the linear mixing method, ABACUS implements Pulay’s direct inversion in iterative subspace (DIIS) method Pulay (1980, 1982) and Broyden’s quasi-Newton approach. Johnson (1988) Fig. 3 illustrates the electron density mixing strategies implemented in ABACUS to guarantee stable SCF convergence.
Magnetic calculations usually exhibit significantly poorer convergence behavior compared to their non-magnetic counterparts, frequently encountering two distinct failure cases, i.e., divergence of SCF or convergence to excited magnetic configurations. These issues are due to the complex coupling between charge and magnetization degrees of freedom when solving the Kohn-Sham equations. To address these challenges, ABACUS implements a mixing scheme that explicitly accounts for spin polarization. For collinear cases, the residual vector $R$ is reformulated to simultaneously track both charge and magnetization density variations
@@
where $n$ determines the history length, $p_{j}$ are Pulay/Broyden coefficients determined according to residual vector (Eq. 14), and the $q$-dependent terms represents a modified Kerker preconditioner. Zhou et al. (2018) The damping parameters $q_{\rho}$ and $q_{m}$ control the suppression of long-wavelength oscillations in charge and magnetization densities, respectively.
ABACUS provides several choices to optimize magnetic calculations, beginning with the implementation of separate preconditioning parameters for charge $q_{\rho}$ and magnetization $q_{m}$ densities to address their distinct physical characteristics and convergence behaviors. By default, the code deactivates magnetic density preconditioning (setting $q_{m}$ = 0) to prioritize numerical stability in most cases, while still providing users the flexibility to enable magnetic preconditioning when enhanced convergence for challenging magnetic systems is required. This approach combines robust default settings with customizable options, allowing the code to automatically handle routine cases while permitting users to fine-tune the mixing parameters for complex magnetic configurations. We compared the magnetic convergence for several magnetic materials using two versions of the ABACUS (v3.4 and v3.5). All computational details are available online. mag (2024) As illustrated in Fig. 4(a), a marked difference in SCF convergence behavior is observed between the two versions. Our findings demonstrate that ABACUS v3.5 achieves significantly enhanced convergence efficiency when compared to the version 3.4 of ABACUS.
@@
Figure 4: (a) Differences in SCF performance score (vide infra) between v3.4 and v3.5 on the selection set of 21 examples. Prior to v3.4, the residual was defined by $\rho_{\uparrow}$ and $\rho_{\downarrow}$, whereas starting from v3.5, the residual is defined by the charge density and the magnetic density, and they are mixed separately. Since some test cases failed to converge in v3.4, we present results with convergence scores rather than directly using convergence steps. The max number of iterations is set to 100 in all calculations, and the SCF convergence threshold scf_thr=1e-6. The SCF score is defined by $|\rm log_{10}(\delta\rho_{\rm last})|\times 10\times\frac{100}{\rm N_{step}}$, where $\delta\rho_{\rm last}$ is the density difference $R(\rho_{in})$ of the last iteration and $\rm N_{step}$ is the number of convergence iterations. (b) Difference in SCF convergence step ($\rm N_{step}$) for the DFT+U calculation with only mixing charge density and both mixing charge density and density matrix. Here, “JT” represents a structure under Jahn–Teller distortion. Halcrow (2013) For the sake of reproduction, we have made public all the details of the calculations in the link, mag (2024) where one can find the complete report and download all input files.
@@
The charge mixing scheme in DFT+U calculations can be further enhanced by an automated U-Ramping procedure that gradually increases the U parameter during SCF iterations. Meredig et al. (2010); Dorado et al. (2013) We performed DFT+U calculations on a series of selected magnetic materials (source files available online mag (2024)) using ABACUS v3.6, with particular focus on comparing the SCF convergence behavior between calculations with and without density matrix mixing. Fig. 4(b) demonstrates the reduction of SCF convergence steps after the density matrix mixing method is adopted.
@@ II.2.2 Smearing Methods
In metallic systems, SCF iterations frequently encounter convergence difficulties stemming from charge density oscillations. These oscillations arise because the electron occupation number drops abruptly to zero at the Fermi energy, with electronic states near the Fermi surface frequently switching between filled and empty electronic states. Furthermore, there exists no variational principle that regulates the convergence of the system’s total energy with resepct to $k$-point sampling. Smearing methods mitigate this issue by introducing a controlled degree of partial occupation for Kohn-Sham orbitals in the vicinity of the Fermi energy, effectively damping numerical instabilities and promoting convergence. A few smearing methods that have been implemented in ABACUS are listed below.
@@
Second, the Fermi-Dirac smearing scheme is intrinsically aligned with finite-temperature density functional theory, Mermin (1965) where the electronic temperature $T$ governs the statistical electron distribution. Crucially, within this framework the smearing width ($k_{B}T$) arises as a natural consequence of thermal excitations in the electronic system, rather than serving as an empirically tuned broadening parameter. The formula is
@@
where $k_{B}$ denotes Boltzmann’s constant. However, in comparison to Gaussian smearing, the drawback of the Fermi-Dirac smearing method lies in its long tail, which necessitates the calculation and storage of a large number of partially occupied states to obtain the charge density.
@@
Fourth, to address possible negative occupancies that may lead to negative total electron density during SCF and numerical instability in the MP method, Marzari et al. introduced the cold-smearing method, Marzari et al. (1999) where the broadening function takes the form of
@@ II.3 Geometry Relaxation
This minimization problem can be solved using various optimization algorithms, including the steepest descent (SD), the conjugate gradient (CG), the Broyden-Fletcher-Goldfarb-Shanno (BFGS), Head and Zerner (1985) and the Fast Inertial Relaxation Engine (FIRE) Bitzek et al. (2006) methods. At each iteration step, the atomic positions are updated according to
@@
where $\Omega$ denotes the cell volume and $\epsilon_{\alpha\beta}$ is the strain tensor components. Next, different geometry relaxation strategies can be employed. For example, one approach to perform cell relaxation consists of two sequential optimization stages. First, optimizing atomic positions with lattice vectors held fixed, followed by adjusting lattice vectors while keeping ionic positions constant. An alternative strategy involves optimization of both atomic positions and lattice vectors at the same time. This simultaneous refinement allows for dynamic coupling between atomic rearrangements and lattice deformation, potentially accelerating convergence to the equilibrium structure. The geometry optimization methods implemented in ABACUS have been applied to study alloys, Liu et al. (2024a) interfaces, Liu, Liu, and Chen (2021) slabs, Liu et al. (2022) and low-dimensional materials, Chen et al. (2022b); Liu and Chen (2025) etc.
@@ II.4 Molecular Dynamics
Central to MD simulations is the potential energy surface (PES), which describes atomic interactions and can be derived through various methods. Quantum mechanical approaches like density functional theory (DFT) can provide high-accuracy PES calculations but are computationally intensive. Alternatively, machine learning potentials Wang et al. (2018); Zeng et al. (2023) offer a balance between accuracy and efficiency by learning the PES from DFT training data, enabling simulations of larger systems or longer timescales. Although classical MD methods enable rapid PES evaluations for systems, their reliance on empirical force fields (such as Lennard-Jones potential Jones and Chapman (1924a, b)) results in lower accuracy compared to the other two approaches.
Once the PES is established, atomic forces are computed via gradients of the total energy with respect to atomic positions, which are then integrated over time to predict atomic trajectories. The velocity Verlet algorithm, Swope et al. (1982) a widely used integration scheme that ensures long-term stability, updates atomic positions $\mathbf{r}(t+\Delta t)$ and velocities $\mathbf{v}(t+\Delta t)$ in discrete time steps $\Delta t$.
MD simulations can be performed under various statistical ensembles, each imposing different constraints to model specific experimental conditions. ABACUS supports the following ensembles. First, the NVE (microcanonical) ensemble conserves particle number $N$, volume $V$, and total energy $E$, simulating isolated systems where energy exchange with the environment is negligible. Second, for simulations requiring constant temperature $T$, the NVT (canonical) ensemble Glenn J. Martyna and Klein (1996) employs thermostats such as velocity rescaling, a straightforward but non-equilibrium approach, or the Nosé-Hoover chain method, Nosé (1984); Hoover (1985); Martyna, Klein, and Tuckerman (1992) which rigorously incorporates auxiliary degrees of freedom to maintain temperature via Hamiltonian dynamics. The Anderson thermostat Andersen (1980) couples the system to a heat bath that imposes the desired temperature to simulate the NVT ensemble. The coupling to a heat bath is represented by a stochastic collision that occasionally acts on randomly selected particles. The Berendsen thermostat Berendsen et al. (1984) rescales the velocities of atoms at each step to reset the temperature of a group of atoms. In addition, the Langevin dynamics Schneider and Stoll (1978) couples the system to a heat bath through stochastic and frictional forces, ensuring that the system reaches and maintains the desired temperature over time, mimicking conditions where the system exchanges energy with its environment. Third, the NPT (isothermal-isobaric) ensemble maintains constant pressure $P$ and temperature $T$ using barostats like Parrinello-Rahman.
Besides the above methods, ABACUS also supports the multi-scale shock technique (MSST), Reed, Fried, and Joannopoulos (2003) enabling the simulation of shock wave propagation in materials, thereby extending the range of accessible timescales for investigating shock phenomena within such non-equilibrium frameworks.
@@
Figure 6: Radial distribution functions (RDFs) $g(r)$ for B systems at a density of 2.46 $\mathrm{g/cm^{3}}$ and the temperatures are (a) 86 eV and (b) 350 eV. Some of the $g(r)$ are derived from Ext-FPMD, Blanchet et al. (2022a) as well as the sDFT calculations employed the PBE Perdew, Burke, and Ernzerhof (1996) and corrKSDT Karasiev, Dufty, and Trickey (2018) XC functionals. The number of B atoms is set to 32 in AIMD simulations. In addition, DPMD denotes the model trained by the traditional DP method, Wang et al. (2018) whereas DPMD-T indicates the TDDP method applied model training as reported by Zhang et al. Zhang et al. (2020b) NN is the number of B atoms in a cell. (Adapted with permission from Matter. Radiat. Extremes 9, 015604 (2024). Copyright 2024 AIP Publishing.)
Chen et al. Chen et al. (2024a) combined stochastic density functional theory (sDFT) with BOMD to explore warm dense matter systems at temperatures ranging from several tens of eV to 1000 eV. They also trained machine-learning-based interatomic models using the first-principles data and employed these models to examine large systems via BOMD simulations. Furthermore, they evaluated the structural and dynamic characteristics, as well as the transport coefficients, of warm dense matter. Fig. 6 shows the radial distribution functions $g(r)$ of warm dense B with a density of 2.46 $\mathrm{g/cm^{3}}$ at 86 and 350 eV, and the sDFT results are in excellent agreement with those obtained from extended first-principles molecular dynamics (Ext-FPMD). Zhang et al. (2016); Blanchet et al. (2022b)
Ma et al. Ma et al. (2024) combine AIMD and finite-temperature orbital-free DFT (FT-OFDFT) with a nonlocal free energy functional XWM to study a variety of warm dense matter systems such as the Si, Al, H, He, and H-He mixtures. The KS-DFT calculations were performed using ABACUS, while the OF-DFT calculations were carried out with ATLAS. Mi et al. (2016) The XWMF functional is expected to be a good choice for the realistic simulations of warm dense matter systems covering a broad range of temperatures and pressures.
@@ II.5 Implicit Solvation Model
We place the solute in a cavity surrounded by a continuum dielectric medium characterized by the relative permittivity of the solvent, as employed in a previous work. Mathew et al. (2014) We describe the dielectric response in terms of the solute’s electron density, considering the polarization of solvents in response to the electronic structures of solute, the effects of cavitation and dispersion, and the reaction of solute system to the presence of the solvent.
@@
We benchmark the accuracy of the implicit solvation implementation by calculating molecular solvation energies $E_{\mathrm{sol}}$, which is defined as the total energy difference between a solvated condition and a vacuum condition in ABACUS and comparing them against VASPsol-calculated values. Mathew et al. (2019) We can see from Fig. 8 that the solvation energies from ABACUS agree well with those produced by the VASPsol package. The minor discrepancies between the solvation energies computed by the two methods may be attributed to the pseudopotential difference. Note that we use projector-augmented-wave (PAW) potentials in VASPsol and the norm-conserving pseudopotentials in ABACUS.
@@
ABACUS also supports adding dipole correction and compensating charge when modeling surfaces, which were utilized in a recent work. Sun et al. (2025) The periodic boundary conditions imposed on the electrostatic potential create an artificial electric field across a slab. By introducing an isolated slab-shaped density distribution $\rho(\mathbf{r})$ that is normal to the $z$-axis, a dipole correction Bengtsson (1999) is added to the bare ionic potential to compensate for the artificial dipole field within the context of periodic supercell calculations.
@@
Modeling a constant-potential electrochemical surface reaction requires the adjustment of electron numbers in a simulation cell. Simultaneously, we must preserve the supercell’s neutrality under the periodic boundary conditions. Thus, a distribution of compensating charge needs to be implemented in the vacuum region of surface models when extra electrons are added to or extracted from the system. The compensating charge implemented in ABACUS follows the methodology developed by Brumme et al. Brumme, Calandra, and Mauri (2014) We assume that the monopole with a total charge of $n_{\mathrm{dop}}$ per unit cell is located at $z_{\mathrm{mono}}$ along the $z$ axis. In this case, the effective potential $V_{\mathrm{mono}}(\mathbf{r})$ is added to the Hamiltonian, and the additional term $E_{\mathrm{mono}}$ is included in the total energy. Since $V_{\mathrm{mono}}(\mathbf{r})$ is independent of the electron density, it is unnecessary to update in a self-consistent manner. We also implemented the correction on the ionic forces induced by the presence of the monopole, allowing for calculating the electronic structure and complete structural relaxation in the field-effect configuration.
@@
Figure 9: Electrostatic energy of an electron along the $z$ direction perpendicular to the Pt surface calculated by Quantum ESPRESSO and ABACUS. Sun et al. (2025) The gray vertical dashed line represents the position of the compensating charge plate, while the blue dashed line corresponds to the dipole correction. A schematic structure plot of the testing interface model is aligned below the average electrostatic energy curves. Gray, red, and white spheres correspond to platinum, oxygen, and hydrogen atoms, respectively.
@@
We output the electrostatic energy of an electron along the $z$ direction perpendicular to the Pt surface for a testing system with an added electron number $N_{e}^{\mathrm{extra}}=0.2$ in the simulation cell. The testing system is a (3×3) Pt (111) surface slab composed of three atomic layers, as illustrated in Fig. 9. The modeled electrode surface contains 27 Pt atoms, with a monolayer of hydrogen coverage and six explicit water molecules. We place a compensating charge plate in the vacuum region above the water layer, and a dipole correction is also included in DFT calculations. Fig. 9 shows that the electrostatic energy results from ABACUS are in perfect agreement with those computed from Quantum ESPRESSO. Giannozzi et al. (2020)